# A Theory of Recommendations

### v2

@@ A.2 Distinct Sender and Receiver Distributions
A.3 Multiple Recommendation Levels
@@
A.4 Multiple Recommendations
@@
First, Lemma 3 implies that infinite learning fully reveals the good and bad products, whereas the posteriors of the controversial product versions are $p_{i}(b,d)=q_{i}/(q_{1}+q_{2})$ for $b,d>0$ Hence, the expected payoff of buying the recommended good with mixed recommendations reads
@@
    If $\lambda>1$, only $i\geq\tilde{i}_{\infty}$ buy the product by Lemma 4. Hence, no type buys a controversial product in the infinite-learning recommendation system if $\tilde{i}_{\infty}\geq 1/2$, which is equivalent to $\sigma\geq\lambda$. This inequality implies $\sigma>1$. Hence, in that case all receivers behave in the same way in both the optimal single-recommendation system (Proposition 5) and the infinite-learning system, buying only objectively good products. Conversely, if $\tilde{i}_{\infty}\leq-1/2$ all types buy a controversial product with the infinite-learning recommendation system. This can be rewritten as $\sigma\leq 1/\lambda$, which implies $\sigma<1$, so that all receivers buy all products except those that are objectively bad and therefore behave in the same way in the optimal single-recommendation system and the infinite-learning system.

### v3

@@ A.3 Distinct Sender and Receiver Distributions
A.4 Multiple Recommendation Levels
@@
A.5 Multiple Recommendations
We now derive some auxiliary results for the case of multiple recommendations, before we prove Proposition 11.
@@ A.5 Multiple Recommendations
First, Lemma A.4 implies that infinite learning fully reveals the good and bad products, whereas the posteriors of the controversial product versions are $p_{i}(b,d)=q_{i}/(q_{1}+q_{2})$ for $b,d>0$ Hence, the expected payoff of buying the recommended good with mixed recommendations reads
@@
    If $\lambda>1$, only $i\geq\tilde{i}_{\infty}$ buy the product by Lemma A.5. Hence, no type buys a controversial product in the infinite-learning recommendation system if $\tilde{i}_{\infty}\geq 1/2$, which is equivalent to $\sigma\geq\lambda$. This inequality implies $\sigma>1$. Hence, in that case all receivers behave in the same way in both the optimal single-recommendation system (Proposition 5) and the infinite-learning system, buying only objectively good products. Conversely, if $\tilde{i}_{\infty}\leq-1/2$ all types buy a controversial product with the infinite-learning recommendation system. This can be rewritten as $\sigma\leq 1/\lambda$, which implies $\sigma<1$, so that all receivers buy all products except those that are objectively bad and therefore behave in the same way in the optimal single-recommendation system and the infinite-learning system.
@@
Appendix B Interactive Plots of $V(R)$ and $D(R)$
In what follows, we describe how to generate interactive plots of $V(R)$ and $D(R)$ under Assumption 2. The numerical analysis is based on a self-contained R script using the shiny and ggplot2 packages. The script allows users to manipulate model parameters in real time and visualize the corresponding plots through a simple web-based interface.
Requirements
  • R (version 4.0 or higher recommended)
  • RStudio (optional but recommended)
The script requires the R packages shiny and ggplot2, both of which are automatically checked for and installed if necessary. No manual installation is required. Beyond that, you only need to download the file recommendations_simulations.R . This single R script contains the entire code for both value and demand plots.
Instructions for Use
  1. 1.
    Open RStudio.
  2. 2.
    Open the file recommendations_simulations.R .
  3. 3.
    Run the script (e.g., select all lines and press Ctrl + Enter or click “Run App”).
  4. 4.
    A Shiny application will launch in your Viewer pane or in a web browser.
  5. 5.
    Use the sliders to manipulate the parameters $\sigma$, $a$ and $Q$.
  6. 6.
    Navigate between the two tabs:
    • V(R): Value of the recommendation system
    • D(R): Demand for the recommended product
Technical Details
  • The $\sigma$ and $a$ sliders operate on a logarithmic scale centered around $1$ to enable easy exploration.
  • The $R$ thresholds are plotted from $0.001$ to $0.999$.
  • The app will dynamically update plots based on the current input values.